rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#5477 — FS#9397 — Urgent maintenance Roubaix <> Lille cable - 24/09 from 10:00pm

Attached to Project— Network
Maintenance
Whole Network
CLOSED
100%
This task is following the incident that we have had yesterday (task #9382) on one of our fibers in Lille. Cutoff due to GC works in Roubaix which has damaged the cable without cutting it off completely.

The workaround set yesterday has allowed us to rapidly and relatively recover the circuits by switching the services on the pairs that haven't been touched. The cable remains therefore very fragile at the level of impact and we have also been notified that setting a high voltage cable a few centimetres below the damaged cover's part would be performed tomorrow.
We have decided to organise an urgent maintenance tonight, to reset the cable and a new cover a few meters of this zone, in order to secure the link and avoid a cutoff during the day. However, it is a provisional reparation and we have to organise another final reparation within few months.

The operation will be performed tonight from 10:00pm and will last from 2 to 3 hours.

Expected impacts:
- The backbone links to London and Paris TH2 will be interrupted but will automatically switch in protection.
- The Euratechno link will be cutoff and the internet access will remain unavailable during the intervention.
- One of the NRA loop tabs will be cutoff but with no impact on the traffic.
- The NRAs in Haubourdin and Canteleu which are in pendulum will be isolated and thus, the access to the subscribers will be interrupted during the maintenance.

Date:  Wednesday, 25 September 2013, 00:44AM
Reason for closing:  Done
Additional comments about closing:  All links are performing.
Comment by OVH - Wednesday, 25 September 2013, 00:36AM

We are starting the works.


Comment by OVH - Wednesday, 25 September 2013, 00:37AM

The damaged cable is cutoff. The new cover is being installed.


Comment by OVH - Wednesday, 25 September 2013, 00:37AM

The cover is set. The new cable welding starts.


Comment by OVH - Wednesday, 25 September 2013, 00:39AM

All fibers are welded. We are doing the checks. The 2 NRAs isolated by the maintenance are online as of around twenty minutes.


Comment by OVH - Wednesday, 25 September 2013, 00:43AM

We are resetting the RAMAN script on Roubaix to London to fix the amplifier taking into account the changes in the characteristics of the fiber.
All other links are performing.